$SPAD/src/input nsfip.input

Timothy Daly December 3, 2016

Abstract

Contents

2

1 License

— license —

--Copyright The Numerical Algorithms Group Limited 1991.

———-

— * —

)set break resume
)spool nsfip.output
)set message test on
)set message auto off
)clear all

———-

Note that the conversions of Results from DoubleFloat to Float will become unnecessary if outputGeneral is extended to apply to DoubleFloat quantities.

— * —

--S 1 of 141
outputGeneral 4
--R
--R                                                                   Type: Void
--E 1

———-

DF here means DoubleFloat. Results converted to Float as outputGeneral not working on DF.

nagExpInt : DF -¿ DF ; — * —

--S 2 of 141 used to work?
nagExpInt(2) :: Float
--R
--R   There are no library operations named nagExpInt
--R      Use HyperDoc Browse or issue
--R                             )what op nagExpInt
--R      to learn if there is any operation containing " nagExpInt " in
--R      its name.
--R
--RDaly Bug
--R   Cannot find a definition or applicable library operation named
--R      nagExpInt with argument type(s)
--R                               PositiveInteger

--R
--R      Perhaps you should use "@" to indicate the required return type,
--R      or "$" to specify which version of the function you need.
--E 2
--       0.0489
--S 3 of 141
nagExpInt(-1) :: Float
--R
--R   There are no library operations named nagExpInt
--R      Use HyperDoc Browse or issue
--R                             )what op nagExpInt
--R      to learn if there is any operation containing " nagExpInt " in
--R      its name.
--R
--RDaly Bug
--R   Cannot find a definition or applicable library operation named
--R      nagExpInt with argument type(s)
--R                                   Integer
--R
--R      Perhaps you should use "@" to indicate the required return type,
--R      or "$" to specify which version of the function you need.
--E 3
--
-- ** ABNORMAL EXIT from NAG Library routine S13AAF: IFAIL =     1
-- ** NAG soft failure - control returned
--
--   Error signalled from user code:
--      An error was detected when calling the NAG Library routine
--      S13AAF. The error number (IFAIL value) is 1, please consult the
--      NAG manual via the Browser for diagnostic information.

———-

nagSinInt : DF -¿ DF ; — * —

--S 4 of 141 used to work?
nagSinInt(0) :: Float
--R
--R   There are no library operations named nagSinInt
--R      Use HyperDoc Browse or issue
--R                             )what op nagSinInt
--R      to learn if there is any operation containing " nagSinInt " in
--R      its name.
--R
--RDaly Bug
--R   Cannot find a definition or applicable library operation named
--R      nagSinInt with argument type(s)
--R                             NonNegativeInteger

--R
--R      Perhaps you should use "@" to indicate the required return type,
--R      or "$" to specify which version of the function you need.
--E 4
--       0.0
--S 5 of 141 used to work?
nagSinInt(0.2) :: Float
--R
--R   There are no library operations named nagSinInt
--R      Use HyperDoc Browse or issue
--R                             )what op nagSinInt
--R      to learn if there is any operation containing " nagSinInt " in
--R      its name.
--R
--RDaly Bug
--R   Cannot find a definition or applicable library operation named
--R      nagSinInt with argument type(s)
--R                                    Float
--R
--R      Perhaps you should use "@" to indicate the required return type,
--R      or "$" to specify which version of the function you need.
--E 5
--       0.1996
--S 6 of 141 used to work?
nagSinInt(0.4) :: Float
--R
--R   There are no library operations named nagSinInt
--R      Use HyperDoc Browse or issue
--R                             )what op nagSinInt
--R      to learn if there is any operation containing " nagSinInt " in
--R      its name.
--R
--RDaly Bug
--R   Cannot find a definition or applicable library operation named
--R      nagSinInt with argument type(s)
--R                                    Float
--R
--R      Perhaps you should use "@" to indicate the required return type,
--R      or "$" to specify which version of the function you need.
--E 6
--       0.3965
--S 7 of 141 used to work?
nagSinInt(0.6) :: Float
--R
--R   There are no library operations named nagSinInt
--R      Use HyperDoc Browse or issue
--R                             )what op nagSinInt

--R      to learn if there is any operation containing " nagSinInt " in
--R      its name.
--R
--RDaly Bug
--R   Cannot find a definition or applicable library operation named
--R      nagSinInt with argument type(s)
--R                                    Float
--R
--R      Perhaps you should use "@" to indicate the required return type,
--R      or "$" to specify which version of the function you need.
--E 7
--       0.5881
--S 8 of 141 used to work?
nagSinInt(0.8) :: Float
--R
--R   There are no library operations named nagSinInt
--R      Use HyperDoc Browse or issue
--R                             )what op nagSinInt
--R      to learn if there is any operation containing " nagSinInt " in
--R      its name.
--R
--RDaly Bug
--R   Cannot find a definition or applicable library operation named
--R      nagSinInt with argument type(s)
--R                                    Float
--R
--R      Perhaps you should use "@" to indicate the required return type,
--R      or "$" to specify which version of the function you need.
--E 8
--       0.7721
--S 9 of 141 used to work?
nagSinInt(1) :: Float
--R
--R   There are no library operations named nagSinInt
--R      Use HyperDoc Browse or issue
--R                             )what op nagSinInt
--R      to learn if there is any operation containing " nagSinInt " in
--R      its name.
--R
--RDaly Bug
--R   Cannot find a definition or applicable library operation named
--R      nagSinInt with argument type(s)
--R                               PositiveInteger
--R
--R      Perhaps you should use "@" to indicate the required return type,
--R      or "$" to specify which version of the function you need.
--E 9
--       0.9461

———-

nagCosInt : DF -¿ DF ; — * —

--S 10 of 141 used to work?
nagCosInt(0.2) :: Float
--R
--R   There are no library operations named nagCosInt
--R      Use HyperDoc Browse or issue
--R                             )what op nagCosInt
--R      to learn if there is any operation containing " nagCosInt " in
--R      its name.
--R
--RDaly Bug
--R   Cannot find a definition or applicable library operation named
--R      nagCosInt with argument type(s)
--R                                    Float
--R
--R      Perhaps you should use "@" to indicate the required return type,
--R      or "$" to specify which version of the function you need.
--E 10
--       - 1.042
--S 11 of 141 used to work?
nagCosInt(0.4) :: Float
--R
--R   There are no library operations named nagCosInt
--R      Use HyperDoc Browse or issue
--R                             )what op nagCosInt
--R      to learn if there is any operation containing " nagCosInt " in
--R      its name.
--R
--RDaly Bug
--R   Cannot find a definition or applicable library operation named
--R      nagCosInt with argument type(s)
--R                                    Float
--R
--R      Perhaps you should use "@" to indicate the required return type,
--R      or "$" to specify which version of the function you need.
--E 11
--       - 0.3788
--S 12 of 141
nagCosInt(0.6) :: Float
--R
--R   There are no library operations named nagCosInt
--R      Use HyperDoc Browse or issue

--R                             )what op nagCosInt
--R      to learn if there is any operation containing " nagCosInt " in
--R      its name.
--R
--RDaly Bug
--R   Cannot find a definition or applicable library operation named
--R      nagCosInt with argument type(s)
--R                                    Float
--R
--R      Perhaps you should use "@" to indicate the required return type,
--R      or "$" to specify which version of the function you need.
--E 12
--       - 0.02227
--S 13 of 141
nagCosInt(0.8) :: Float
--R
--R   There are no library operations named nagCosInt
--R      Use HyperDoc Browse or issue
--R                             )what op nagCosInt
--R      to learn if there is any operation containing " nagCosInt " in
--R      its name.
--R
--RDaly Bug
--R   Cannot find a definition or applicable library operation named
--R      nagCosInt with argument type(s)
--R                                    Float
--R
--R      Perhaps you should use "@" to indicate the required return type,
--R      or "$" to specify which version of the function you need.
--E 13
--       0.1983
--S 14 of 141
nagCosInt(1) :: Float
--R
--R   There are no library operations named nagCosInt
--R      Use HyperDoc Browse or issue
--R                             )what op nagCosInt
--R      to learn if there is any operation containing " nagCosInt " in
--R      its name.
--R
--RDaly Bug
--R   Cannot find a definition or applicable library operation named
--R      nagCosInt with argument type(s)
--R                               PositiveInteger
--R
--R      Perhaps you should use "@" to indicate the required return type,
--R      or "$" to specify which version of the function you need.
--E 14

--       0.3374

———-

nagIncompleteGammaP : (DF, DF) -¿ DF ; (to machine precision)

— * —

--S 15 of 141
nagIncompleteGammaP(2,3) :: Float
--R
--R   There are no library operations named nagIncompleteGammaP
--R      Use HyperDoc Browse or issue
--R                        )what op nagIncompleteGammaP
--R      to learn if there is any operation containing "
--R      nagIncompleteGammaP " in its name.
--R
--RDaly Bug
--R   Cannot find a definition or applicable library operation named
--R      nagIncompleteGammaP with argument type(s)
--R                               PositiveInteger
--R                               PositiveInteger
--R
--R      Perhaps you should use "@" to indicate the required return type,
--R      or "$" to specify which version of the function you need.
--E 15
--       0.8009
--S 16 of 141
nagIncompleteGammaP(7,1) :: Float
--R
--R   There are no library operations named nagIncompleteGammaP
--R      Use HyperDoc Browse or issue
--R                        )what op nagIncompleteGammaP
--R      to learn if there is any operation containing "
--R      nagIncompleteGammaP " in its name.
--R
--RDaly Bug
--R   Cannot find a definition or applicable library operation named
--R      nagIncompleteGammaP with argument type(s)
--R                               PositiveInteger
--R                               PositiveInteger
--R
--R      Perhaps you should use "@" to indicate the required return type,
--R      or "$" to specify which version of the function you need.
--E 16
--       0.00008324
--S 17 of 141
nagIncompleteGammaP(0.5,99) :: Float

--R
--R   There are no library operations named nagIncompleteGammaP
--R      Use HyperDoc Browse or issue
--R                        )what op nagIncompleteGammaP
--R      to learn if there is any operation containing "
--R      nagIncompleteGammaP " in its name.
--R
--RDaly Bug
--R   Cannot find a definition or applicable library operation named
--R      nagIncompleteGammaP with argument type(s)
--R                                    Float
--R                               PositiveInteger
--R
--R      Perhaps you should use "@" to indicate the required return type,
--R      or "$" to specify which version of the function you need.
--E 17
--       1.0
--S 18 of 141
nagIncompleteGammaP(20,21) :: Float
--R
--R   There are no library operations named nagIncompleteGammaP
--R      Use HyperDoc Browse or issue
--R                        )what op nagIncompleteGammaP
--R      to learn if there is any operation containing "
--R      nagIncompleteGammaP " in its name.
--R
--RDaly Bug
--R   Cannot find a definition or applicable library operation named
--R      nagIncompleteGammaP with argument type(s)
--R                               PositiveInteger
--R                               PositiveInteger
--R
--R      Perhaps you should use "@" to indicate the required return type,
--R      or "$" to specify which version of the function you need.
--E 18
--       0.6157
--S 19 of 141
nagIncompleteGammaP(21,20) :: Float
--R
--R   There are no library operations named nagIncompleteGammaP
--R      Use HyperDoc Browse or issue
--R                        )what op nagIncompleteGammaP
--R      to learn if there is any operation containing "
--R      nagIncompleteGammaP " in its name.
--R
--RDaly Bug
--R   Cannot find a definition or applicable library operation named
--R      nagIncompleteGammaP with argument type(s)

--R                               PositiveInteger
--R                               PositiveInteger
--R
--R      Perhaps you should use "@" to indicate the required return type,
--R      or "$" to specify which version of the function you need.
--E 19
--       0.4409

———-

nagIncompleteGammaP : (DF, DF, DF) -¿ DF ; (to specified precision) — * —

--S 20 of 141
nagIncompleteGammaP(7,1,0.1) :: Float
--R
--R   There are no library operations named nagIncompleteGammaP
--R      Use HyperDoc Browse or issue
--R                        )what op nagIncompleteGammaP
--R      to learn if there is any operation containing "
--R      nagIncompleteGammaP " in its name.
--R
--RDaly Bug
--R   Cannot find a definition or applicable library operation named
--R      nagIncompleteGammaP with argument type(s)
--R                               PositiveInteger
--R                               PositiveInteger
--R                                    Float
--R
--R      Perhaps you should use "@" to indicate the required return type,
--R      or "$" to specify which version of the function you need.
--E 20
--       0.00008313

———-

nagIncompleteGammaQ : (DF, DF) -¿ DF ; (to machine precision) — * —

--S 21 of 141
nagIncompleteGammaQ(2,3) :: Float
--R
--R   There are no library operations named nagIncompleteGammaQ
--R      Use HyperDoc Browse or issue
--R                        )what op nagIncompleteGammaQ
--R      to learn if there is any operation containing "
--R      nagIncompleteGammaQ " in its name.
--R
--RDaly Bug

--R   Cannot find a definition or applicable library operation named
--R      nagIncompleteGammaQ with argument type(s)
--R                               PositiveInteger
--R                               PositiveInteger
--R
--R      Perhaps you should use "@" to indicate the required return type,
--R      or "$" to specify which version of the function you need.
--E 21
--       0.1991
--S 22 of 141
nagIncompleteGammaQ(7,1) :: Float
--R
--R   There are no library operations named nagIncompleteGammaQ
--R      Use HyperDoc Browse or issue
--R                        )what op nagIncompleteGammaQ
--R      to learn if there is any operation containing "
--R      nagIncompleteGammaQ " in its name.
--R
--RDaly Bug
--R   Cannot find a definition or applicable library operation named
--R      nagIncompleteGammaQ with argument type(s)
--R                               PositiveInteger
--R                               PositiveInteger
--R
--R      Perhaps you should use "@" to indicate the required return type,
--R      or "$" to specify which version of the function you need.
--E 22
--       0.9999
--S 23 of 141
nagIncompleteGammaQ(0.5,99) :: Float
--R
--R   There are no library operations named nagIncompleteGammaQ
--R      Use HyperDoc Browse or issue
--R                        )what op nagIncompleteGammaQ
--R      to learn if there is any operation containing "
--R      nagIncompleteGammaQ " in its name.
--R
--RDaly Bug
--R   Cannot find a definition or applicable library operation named
--R      nagIncompleteGammaQ with argument type(s)
--R                                    Float
--R                               PositiveInteger
--R
--R      Perhaps you should use "@" to indicate the required return type,
--R      or "$" to specify which version of the function you need.
--E 23
--       0.5705 E -44

--S 24 of 141
nagIncompleteGammaQ(20,21) :: Float
--R
--R   There are no library operations named nagIncompleteGammaQ
--R      Use HyperDoc Browse or issue
--R                        )what op nagIncompleteGammaQ
--R      to learn if there is any operation containing "
--R      nagIncompleteGammaQ " in its name.
--R
--RDaly Bug
--R   Cannot find a definition or applicable library operation named
--R      nagIncompleteGammaQ with argument type(s)
--R                               PositiveInteger
--R                               PositiveInteger
--R
--R      Perhaps you should use "@" to indicate the required return type,
--R      or "$" to specify which version of the function you need.
--E 24
--       0.3843
--S 25 of 141
nagIncompleteGammaQ(21,20) :: Float
--R
--R   There are no library operations named nagIncompleteGammaQ
--R      Use HyperDoc Browse or issue
--R                        )what op nagIncompleteGammaQ
--R      to learn if there is any operation containing "
--R      nagIncompleteGammaQ " in its name.
--R
--RDaly Bug
--R   Cannot find a definition or applicable library operation named
--R      nagIncompleteGammaQ with argument type(s)
--R                               PositiveInteger
--R                               PositiveInteger
--R
--R      Perhaps you should use "@" to indicate the required return type,
--R      or "$" to specify which version of the function you need.
--E 25
--       0.5591
--S 26 of 141
nagIncompleteGammaQ(25,14) :: Float
--R
--R   There are no library operations named nagIncompleteGammaQ
--R      Use HyperDoc Browse or issue
--R                        )what op nagIncompleteGammaQ
--R      to learn if there is any operation containing "
--R      nagIncompleteGammaQ " in its name.
--R
--RDaly Bug

--R   Cannot find a definition or applicable library operation named
--R      nagIncompleteGammaQ with argument type(s)
--R                               PositiveInteger
--R                               PositiveInteger
--R
--R      Perhaps you should use "@" to indicate the required return type,
--R      or "$" to specify which version of the function you need.
--E 26
--       0.995

———-

nagIncompleteGammaQ : (DF, DF, DF) -¿ DF ; (to specified precision) — * —

--S 27 of 141
nagIncompleteGammaQ(25,14,0.1) :: Float
--R
--R   There are no library operations named nagIncompleteGammaQ
--R      Use HyperDoc Browse or issue
--R                        )what op nagIncompleteGammaQ
--R      to learn if there is any operation containing "
--R      nagIncompleteGammaQ " in its name.
--R
--RDaly Bug
--R   Cannot find a definition or applicable library operation named
--R      nagIncompleteGammaQ with argument type(s)
--R                               PositiveInteger
--R                               PositiveInteger
--R                                    Float
--R
--R      Perhaps you should use "@" to indicate the required return type,
--R      or "$" to specify which version of the function you need.
--E 27
--       0.9953

———-

nagErf : DF -¿ DF ; — * —

--S 28 of 141
nagErf(-6) :: Float
--R
--R   There are no library operations named nagErf
--R      Use HyperDoc Browse or issue
--R                               )what op nagErf
--R      to learn if there is any operation containing " nagErf " in its
--R      name.

--R
--RDaly Bug
--R   Cannot find a definition or applicable library operation named
--R      nagErf with argument type(s)
--R                                   Integer
--R
--R      Perhaps you should use "@" to indicate the required return type,
--R      or "$" to specify which version of the function you need.
--E 28
--       - 1.0
--S 29 of 141
nagErf(-4.5) :: Float
--R
--R   There are no library operations named nagErf
--R      Use HyperDoc Browse or issue
--R                               )what op nagErf
--R      to learn if there is any operation containing " nagErf " in its
--R      name.
--R
--RDaly Bug
--R   Cannot find a definition or applicable library operation named
--R      nagErf with argument type(s)
--R                                    Float
--R
--R      Perhaps you should use "@" to indicate the required return type,
--R      or "$" to specify which version of the function you need.
--E 29
--       - 1.0
--S 30 of 141
nagErf(-1) :: Float
--R
--R   There are no library operations named nagErf
--R      Use HyperDoc Browse or issue
--R                               )what op nagErf
--R      to learn if there is any operation containing " nagErf " in its
--R      name.
--R
--RDaly Bug
--R   Cannot find a definition or applicable library operation named
--R      nagErf with argument type(s)
--R                                   Integer
--R
--R      Perhaps you should use "@" to indicate the required return type,
--R      or "$" to specify which version of the function you need.
--E 30
--       - 0.8427
--S 31 of 141

nagErf(1) :: Float
--R
--R   There are no library operations named nagErf
--R      Use HyperDoc Browse or issue
--R                               )what op nagErf
--R      to learn if there is any operation containing " nagErf " in its
--R      name.
--R
--RDaly Bug
--R   Cannot find a definition or applicable library operation named
--R      nagErf with argument type(s)
--R                               PositiveInteger
--R
--R      Perhaps you should use "@" to indicate the required return type,
--R      or "$" to specify which version of the function you need.
--E 31
--       0.8427
--S 32 of 141
nagErf(4.5) :: Float
--R
--R   There are no library operations named nagErf
--R      Use HyperDoc Browse or issue
--R                               )what op nagErf
--R      to learn if there is any operation containing " nagErf " in its
--R      name.
--R
--RDaly Bug
--R   Cannot find a definition or applicable library operation named
--R      nagErf with argument type(s)
--R                                    Float
--R
--R      Perhaps you should use "@" to indicate the required return type,
--R      or "$" to specify which version of the function you need.
--E 32
--       1.0
--S 33 of 141
nagErf(6) :: Float
--R
--R   There are no library operations named nagErf
--R      Use HyperDoc Browse or issue
--R                               )what op nagErf
--R      to learn if there is any operation containing " nagErf " in its
--R      name.
--R
--RDaly Bug
--R   Cannot find a definition or applicable library operation named
--R      nagErf with argument type(s)
--R                               PositiveInteger

--R
--R      Perhaps you should use "@" to indicate the required return type,
--R      or "$" to specify which version of the function you need.
--E 33
--       1.0

———-

nagErfC : DF -¿ DF ; — * —

--S 34 of 141
nagErfC(-10) :: Float
--R
--R   There are no library operations named nagErfC
--R      Use HyperDoc Browse or issue
--R                              )what op nagErfC
--R      to learn if there is any operation containing " nagErfC " in its
--R      name.
--R
--RDaly Bug
--R   Cannot find a definition or applicable library operation named
--R      nagErfC with argument type(s)
--R                                   Integer
--R
--R      Perhaps you should use "@" to indicate the required return type,
--R      or "$" to specify which version of the function you need.
--E 34
--       2.0
--S 35 of 141
nagErfC(-1) :: Float
--R
--R   There are no library operations named nagErfC
--R      Use HyperDoc Browse or issue
--R                              )what op nagErfC
--R      to learn if there is any operation containing " nagErfC " in its
--R      name.
--R
--RDaly Bug
--R   Cannot find a definition or applicable library operation named
--R      nagErfC with argument type(s)
--R                                   Integer
--R
--R      Perhaps you should use "@" to indicate the required return type,
--R      or "$" to specify which version of the function you need.
--E 35
--       1.843

--S 36 of 141
nagErfC(0) :: Float
--R
--R   There are no library operations named nagErfC
--R      Use HyperDoc Browse or issue
--R                              )what op nagErfC
--R      to learn if there is any operation containing " nagErfC " in its
--R      name.
--R
--RDaly Bug
--R   Cannot find a definition or applicable library operation named
--R      nagErfC with argument type(s)
--R                             NonNegativeInteger
--R
--R      Perhaps you should use "@" to indicate the required return type,
--R      or "$" to specify which version of the function you need.
--E 36
--       1.0
--S 37 of 141
nagErfC(1) :: Float
--R
--R   There are no library operations named nagErfC
--R      Use HyperDoc Browse or issue
--R                              )what op nagErfC
--R      to learn if there is any operation containing " nagErfC " in its
--R      name.
--R
--RDaly Bug
--R   Cannot find a definition or applicable library operation named
--R      nagErfC with argument type(s)
--R                               PositiveInteger
--R
--R      Perhaps you should use "@" to indicate the required return type,
--R      or "$" to specify which version of the function you need.
--E 37
--       0.1573
--S 38 of 141
nagErfC(15) :: Float
--R
--R   There are no library operations named nagErfC
--R      Use HyperDoc Browse or issue
--R                              )what op nagErfC
--R      to learn if there is any operation containing " nagErfC " in its
--R      name.
--R
--RDaly Bug
--R   Cannot find a definition or applicable library operation named
--R      nagErfC with argument type(s)

--R                               PositiveInteger
--R
--R      Perhaps you should use "@" to indicate the required return type,
--R      or "$" to specify which version of the function you need.
--E 38
--       0.7213 E -99

———-

nagDAiryAi : DF -¿ DF ; — * —

--S 39 of 141
nagDAiryAi(-10) :: Float
--R
--R   There are no library operations named nagDAiryAi
--R      Use HyperDoc Browse or issue
--R                             )what op nagDAiryAi
--R      to learn if there is any operation containing " nagDAiryAi " in
--R      its name.
--R
--RDaly Bug
--R   Cannot find a definition or applicable library operation named
--R      nagDAiryAi with argument type(s)
--R                                   Integer
--R
--R      Perhaps you should use "@" to indicate the required return type,
--R      or "$" to specify which version of the function you need.
--E 39
--       0.9963
--S 40 of 141
nagDAiryAi(-1) :: Float
--R
--R   There are no library operations named nagDAiryAi
--R      Use HyperDoc Browse or issue
--R                             )what op nagDAiryAi
--R      to learn if there is any operation containing " nagDAiryAi " in
--R      its name.
--R
--RDaly Bug
--R   Cannot find a definition or applicable library operation named
--R      nagDAiryAi with argument type(s)
--R                                   Integer
--R
--R      Perhaps you should use "@" to indicate the required return type,
--R      or "$" to specify which version of the function you need.
--E 40
--       - 0.01016

--S 41 of 141
nagDAiryAi(0) :: Float
--R
--R   There are no library operations named nagDAiryAi
--R      Use HyperDoc Browse or issue
--R                             )what op nagDAiryAi
--R      to learn if there is any operation containing " nagDAiryAi " in
--R      its name.
--R
--RDaly Bug
--R   Cannot find a definition or applicable library operation named
--R      nagDAiryAi with argument type(s)
--R                             NonNegativeInteger
--R
--R      Perhaps you should use "@" to indicate the required return type,
--R      or "$" to specify which version of the function you need.
--E 41
--       - 0.2588
--S 42 of 141
nagDAiryAi(1) :: Float
--R
--R   There are no library operations named nagDAiryAi
--R      Use HyperDoc Browse or issue
--R                             )what op nagDAiryAi
--R      to learn if there is any operation containing " nagDAiryAi " in
--R      its name.
--R
--RDaly Bug
--R   Cannot find a definition or applicable library operation named
--R      nagDAiryAi with argument type(s)
--R                               PositiveInteger
--R
--R      Perhaps you should use "@" to indicate the required return type,
--R      or "$" to specify which version of the function you need.
--E 42
--       - 0.1591
--S 43 of 141
nagDAiryAi(5) :: Float
--R
--R   There are no library operations named nagDAiryAi
--R      Use HyperDoc Browse or issue
--R                             )what op nagDAiryAi
--R      to learn if there is any operation containing " nagDAiryAi " in
--R      its name.
--R
--RDaly Bug
--R   Cannot find a definition or applicable library operation named
--R      nagDAiryAi with argument type(s)

--R                               PositiveInteger
--R
--R      Perhaps you should use "@" to indicate the required return type,
--R      or "$" to specify which version of the function you need.
--E 43
--       - 0.0002474
--S 44 of 141
nagDAiryAi(10) :: Float
--R
--R   There are no library operations named nagDAiryAi
--R      Use HyperDoc Browse or issue
--R                             )what op nagDAiryAi
--R      to learn if there is any operation containing " nagDAiryAi " in
--R      its name.
--R
--RDaly Bug
--R   Cannot find a definition or applicable library operation named
--R      nagDAiryAi with argument type(s)
--R                               PositiveInteger
--R
--R      Perhaps you should use "@" to indicate the required return type,
--R      or "$" to specify which version of the function you need.
--E 44
--       - 0.3521 E -9
--S 45 of 141
nagDAiryAi(20) :: Float
--R
--R   There are no library operations named nagDAiryAi
--R      Use HyperDoc Browse or issue
--R                             )what op nagDAiryAi
--R      to learn if there is any operation containing " nagDAiryAi " in
--R      its name.
--R
--RDaly Bug
--R   Cannot find a definition or applicable library operation named
--R      nagDAiryAi with argument type(s)
--R                               PositiveInteger
--R
--R      Perhaps you should use "@" to indicate the required return type,
--R      or "$" to specify which version of the function you need.
--E 45
--       - 0.7586 E -26

———-

nagDAiryAi : CDF -¿ CDF ; — * —

--S 46 of 141

--RnagDAiryAi(0.3+0.4*%i) :: Complex Float
--R
--R   There are no library operations named nagDAiryAi
--R      Use HyperDoc Browse or issue
--R                             )what op nagDAiryAi
--R      to learn if there is any operation containing " nagDAiryAi " in
--R      its name.
--R
--RDaly Bug
--R   Cannot find a definition or applicable library operation named
--R      nagDAiryAi with argument type(s)
--R                                Complex Float
--R
--R      Perhaps you should use "@" to indicate the required return type,
--R      or "$" to specify which version of the function you need.
--E 46
--       - 0.2612 + 0.03848 %i

———-

nagDAiryBi : DF -¿ DF ; — * —

--S 47 of 141
nagDAiryBi(-10) :: Float
--R
--R   There are no library operations named nagDAiryBi
--R      Use HyperDoc Browse or issue
--R                             )what op nagDAiryBi
--R      to learn if there is any operation containing " nagDAiryBi " in
--R      its name.
--R
--RDaly Bug
--R   Cannot find a definition or applicable library operation named
--R      nagDAiryBi with argument type(s)
--R                                   Integer
--R
--R      Perhaps you should use "@" to indicate the required return type,
--R      or "$" to specify which version of the function you need.
--E 47
--       0.1194
--S 48 of 141
nagDAiryBi(-1) :: Float
--R
--R   There are no library operations named nagDAiryBi
--R      Use HyperDoc Browse or issue
--R                             )what op nagDAiryBi
--R      to learn if there is any operation containing " nagDAiryBi " in
--R      its name.

--R
--RDaly Bug
--R   Cannot find a definition or applicable library operation named
--R      nagDAiryBi with argument type(s)
--R                                   Integer
--R
--R      Perhaps you should use "@" to indicate the required return type,
--R      or "$" to specify which version of the function you need.
--E 48
--       0.5924
--S 49 of 141
nagDAiryBi(0) :: Float
--R
--R   There are no library operations named nagDAiryBi
--R      Use HyperDoc Browse or issue
--R                             )what op nagDAiryBi
--R      to learn if there is any operation containing " nagDAiryBi " in
--R      its name.
--R
--RDaly Bug
--R   Cannot find a definition or applicable library operation named
--R      nagDAiryBi with argument type(s)
--R                             NonNegativeInteger
--R
--R      Perhaps you should use "@" to indicate the required return type,
--R      or "$" to specify which version of the function you need.
--E 49
--       0.4483
--S 50 of 141
nagDAiryBi(1) :: Float
--R
--R   There are no library operations named nagDAiryBi
--R      Use HyperDoc Browse or issue
--R                             )what op nagDAiryBi
--R      to learn if there is any operation containing " nagDAiryBi " in
--R      its name.
--R
--RDaly Bug
--R   Cannot find a definition or applicable library operation named
--R      nagDAiryBi with argument type(s)
--R                               PositiveInteger
--R
--R      Perhaps you should use "@" to indicate the required return type,
--R      or "$" to specify which version of the function you need.
--E 50
--       0.9324
--S 51 of 141

nagDAiryBi(5) :: Float
--R
--R   There are no library operations named nagDAiryBi
--R      Use HyperDoc Browse or issue
--R                             )what op nagDAiryBi
--R      to learn if there is any operation containing " nagDAiryBi " in
--R      its name.
--R
--RDaly Bug
--R   Cannot find a definition or applicable library operation named
--R      nagDAiryBi with argument type(s)
--R                               PositiveInteger
--R
--R      Perhaps you should use "@" to indicate the required return type,
--R      or "$" to specify which version of the function you need.
--E 51
--       1436.0
--S 52 of 141
nagDAiryBi(10) :: Float
--R
--R   There are no library operations named nagDAiryBi
--R      Use HyperDoc Browse or issue
--R                             )what op nagDAiryBi
--R      to learn if there is any operation containing " nagDAiryBi " in
--R      its name.
--R
--RDaly Bug
--R   Cannot find a definition or applicable library operation named
--R      nagDAiryBi with argument type(s)
--R                               PositiveInteger
--R
--R      Perhaps you should use "@" to indicate the required return type,
--R      or "$" to specify which version of the function you need.
--E 52
--       0.1429 E 10
--S 53 of 141
nagDAiryBi(20) :: Float
--R
--R   There are no library operations named nagDAiryBi
--R      Use HyperDoc Browse or issue
--R                             )what op nagDAiryBi
--R      to learn if there is any operation containing " nagDAiryBi " in
--R      its name.
--R
--RDaly Bug
--R   Cannot find a definition or applicable library operation named
--R      nagDAiryBi with argument type(s)
--R                               PositiveInteger

--R
--R      Perhaps you should use "@" to indicate the required return type,
--R      or "$" to specify which version of the function you need.
--E 53
--       0.9382 E 26

———-

nagDAiryBi : CDF -¿ CDF ; — * —

--S 54 of 141
nagDAiryBi(0.3+0.4*%i) :: Complex Float
--R
--R   There are no library operations named nagDAiryBi
--R      Use HyperDoc Browse or issue
--R                             )what op nagDAiryBi
--R      to learn if there is any operation containing " nagDAiryBi " in
--R      its name.
--R
--RDaly Bug
--R   Cannot find a definition or applicable library operation named
--R      nagDAiryBi with argument type(s)
--R                               Complex(Float)
--R
--R      Perhaps you should use "@" to indicate the required return type,
--R      or "$" to specify which version of the function you need.
--E 54
--       0.4093 + 0.07966 %i

———-

nagScaledDAiryAi : CDF -¿ CDF ; — * —

--S 55 of 141
nagScaledDAiryAi(0.3+0.4*%i) :: Complex Float
--R
--R   There are no library operations named nagScaledDAiryAi
--R      Use HyperDoc Browse or issue
--R                          )what op nagScaledDAiryAi
--R      to learn if there is any operation containing " nagScaledDAiryAi
--R      " in its name.
--R
--RDaly Bug
--R   Cannot find a definition or applicable library operation named
--R      nagScaledDAiryAi with argument type(s)
--R                               Complex(Float)
--R
--R      Perhaps you should use "@" to indicate the required return type,

--R      or "$" to specify which version of the function you need.
--E 55
--       - 0.2744 - 0.02356 %i

———-

nagScaledDAiryBi : CDF -¿ CDF ; — * —

--S 56 of 141
nagScaledDAiryBi(0.3+0.4*%i) :: Complex Float
--R
--R   There are no library operations named nagScaledDAiryBi
--R      Use HyperDoc Browse or issue
--R                          )what op nagScaledDAiryBi
--R      to learn if there is any operation containing " nagScaledDAiryBi
--R      " in its name.
--R
--RDaly Bug
--R   Cannot find a definition or applicable library operation named
--R      nagScaledDAiryBi with argument type(s)
--R                               Complex(Float)
--R
--R      Perhaps you should use "@" to indicate the required return type,
--R      or "$" to specify which version of the function you need.
--E 56
--       0.3924 + 0.07638 %i

———-

nagHankelH1 : (DF, CDF, Int) -¿ List CDF ; — * —

--S 57 of 141
nagHankelH1(0,0.3+0.4*%i,2) :: Matrix Complex Float
--R
--R   There are no library operations named nagHankelH1
--R      Use HyperDoc Browse or issue
--R                            )what op nagHankelH1
--R      to learn if there is any operation containing " nagHankelH1 " in
--R      its name.
--R
--RDaly Bug
--R   Cannot find a definition or applicable library operation named
--R      nagHankelH1 with argument type(s)
--R                             NonNegativeInteger
--R                               Complex(Float)
--R                               PositiveInteger
--R
--R      Perhaps you should use "@" to indicate the required return type,
--R      or "$" to specify which version of the function you need.

--E 57
--       [0.3466 - 0.5588 %i  - 0.7912 - 0.8178 %i]
--S 58 of 141
nagHankelH1(2.3,2,2) :: Matrix Complex Float
--R
--R   There are no library operations named nagHankelH1
--R      Use HyperDoc Browse or issue
--R                            )what op nagHankelH1
--R      to learn if there is any operation containing " nagHankelH1 " in
--R      its name.
--R
--RDaly Bug
--R   Cannot find a definition or applicable library operation named
--R      nagHankelH1 with argument type(s)
--R                                    Float
--R                               PositiveInteger
--R                               PositiveInteger
--R
--R      Perhaps you should use "@" to indicate the required return type,
--R      or "$" to specify which version of the function you need.
--E 58
--       [0.2721 - 0.7398 %i  0.08902 - 1.412 %i]
--S 59 of 141
nagHankelH1(2.12,-1,2) :: Matrix Complex Float
--R
--R   There are no library operations named nagHankelH1
--R      Use HyperDoc Browse or issue
--R                            )what op nagHankelH1
--R      to learn if there is any operation containing " nagHankelH1 " in
--R      its name.
--R
--RDaly Bug
--R   Cannot find a definition or applicable library operation named
--R      nagHankelH1 with argument type(s)
--R                                    Float
--R                                   Integer
--R                               PositiveInteger
--R
--R      Perhaps you should use "@" to indicate the required return type,
--R      or "$" to specify which version of the function you need.
--E 59
--       [- 0.7722 - 1.693 %i  2.601 + 6.527 %i]

———-

nagHankelH2 : (DF, CDF, Int) -¿ List CDF ; — * —

--S 60 of 141
nagHankelH2(6,3.1-1.6*%i,2) :: Matrix Complex Float
--R
--R   There are no library operations named nagHankelH2
--R      Use HyperDoc Browse or issue
--R                            )what op nagHankelH2
--R      to learn if there is any operation containing " nagHankelH2 " in
--R      its name.
--R
--RDaly Bug
--R   Cannot find a definition or applicable library operation named
--R      nagHankelH2 with argument type(s)
--R                               PositiveInteger
--R                               Complex(Float)
--R                               PositiveInteger
--R
--R      Perhaps you should use "@" to indicate the required return type,
--R      or "$" to specify which version of the function you need.
--E 60
--       [- 1.371 - 1.28 %i  - 1.491 - 5.993 %i]

———-

nagScaledHankelH1 : (DF, CDF, Int) -¿ List CDF ; — * —

--S 61 of 141
nagScaledHankelH1(0,0.3+0.4*%i,2) :: Matrix Complex Float
--R
--R   There are no library operations named nagScaledHankelH1
--R      Use HyperDoc Browse or issue
--R                         )what op nagScaledHankelH1
--R      to learn if there is any operation containing " nagScaledHankelH1
--R      " in its name.
--R
--RDaly Bug
--R   Cannot find a definition or applicable library operation named
--R      nagScaledHankelH1 with argument type(s)
--R                             NonNegativeInteger
--R                               Complex(Float)
--R                               PositiveInteger
--R
--R      Perhaps you should use "@" to indicate the required return type,
--R      or "$" to specify which version of the function you need.
--E 61
--       [0.2477 - 0.9492 %i  - 1.488 - 0.8166 %i]

———-

nagScaledHankelH2 : (DF, CDF, Int) -¿ List CDF ; — * —

--S 62 of 141
nagScaledHankelH2(6,3.1-1.6*%i,2) :: Matrix Complex Float
--R
--R   There are no library operations named nagScaledHankelH2
--R      Use HyperDoc Browse or issue
--R                         )what op nagScaledHankelH2
--R      to learn if there is any operation containing " nagScaledHankelH2
--R      " in its name.
--R
--RDaly Bug
--R   Cannot find a definition or applicable library operation named
--R      nagScaledHankelH2 with argument type(s)
--R                               PositiveInteger
--R                               Complex(Float)
--R                               PositiveInteger
--R
--R      Perhaps you should use "@" to indicate the required return type,
--R      or "$" to specify which version of the function you need.
--E 62
--       [7.05 + 6.052 %i  8.614 + 29.35 %i]

———-

nagKelvinBer : DF -¿ DF ; — * —

--S 63 of 141
nagKelvinBer(0.1) :: Float
--R
--R   There are no library operations named nagKelvinBer
--R      Use HyperDoc Browse or issue
--R                            )what op nagKelvinBer
--R      to learn if there is any operation containing " nagKelvinBer " in
--R      its name.
--R
--RDaly Bug
--R   Cannot find a definition or applicable library operation named
--R      nagKelvinBer with argument type(s)
--R                                    Float
--R
--R      Perhaps you should use "@" to indicate the required return type,
--R      or "$" to specify which version of the function you need.
--E 63
--       1.0
--S 64 of 141
nagKelvinBer(1) :: Float
--R

--R   There are no library operations named nagKelvinBer
--R      Use HyperDoc Browse or issue
--R                            )what op nagKelvinBer
--R      to learn if there is any operation containing " nagKelvinBer " in
--R      its name.
--R
--RDaly Bug
--R   Cannot find a definition or applicable library operation named
--R      nagKelvinBer with argument type(s)
--R                               PositiveInteger
--R
--R      Perhaps you should use "@" to indicate the required return type,
--R      or "$" to specify which version of the function you need.
--E 64
--       0.9844
--S 65 of 141
nagKelvinBer(2.5) :: Float
--R
--R   There are no library operations named nagKelvinBer
--R      Use HyperDoc Browse or issue
--R                            )what op nagKelvinBer
--R      to learn if there is any operation containing " nagKelvinBer " in
--R      its name.
--R
--RDaly Bug
--R   Cannot find a definition or applicable library operation named
--R      nagKelvinBer with argument type(s)
--R                                    Float
--R
--R      Perhaps you should use "@" to indicate the required return type,
--R      or "$" to specify which version of the function you need.
--E 65
--       0.4
--S 66 of 141
nagKelvinBer(5) :: Float
--R
--R   There are no library operations named nagKelvinBer
--R      Use HyperDoc Browse or issue
--R                            )what op nagKelvinBer
--R      to learn if there is any operation containing " nagKelvinBer " in
--R      its name.
--R
--RDaly Bug
--R   Cannot find a definition or applicable library operation named
--R      nagKelvinBer with argument type(s)
--R                               PositiveInteger
--R
--R      Perhaps you should use "@" to indicate the required return type,

--R      or "$" to specify which version of the function you need.
--E 66
--       - 6.23
--S 67 of 141
nagKelvinBer(10) :: Float
--R
--R   There are no library operations named nagKelvinBer
--R      Use HyperDoc Browse or issue
--R                            )what op nagKelvinBer
--R      to learn if there is any operation containing " nagKelvinBer " in
--R      its name.
--R
--RDaly Bug
--R   Cannot find a definition or applicable library operation named
--R      nagKelvinBer with argument type(s)
--R                               PositiveInteger
--R
--R      Perhaps you should use "@" to indicate the required return type,
--R      or "$" to specify which version of the function you need.
--E 67
--       138.8
--S 68 of 141
nagKelvinBer(15) :: Float
--R
--R   There are no library operations named nagKelvinBer
--R      Use HyperDoc Browse or issue
--R                            )what op nagKelvinBer
--R      to learn if there is any operation containing " nagKelvinBer " in
--R      its name.
--R
--RDaly Bug
--R   Cannot find a definition or applicable library operation named
--R      nagKelvinBer with argument type(s)
--R                               PositiveInteger
--R
--R      Perhaps you should use "@" to indicate the required return type,
--R      or "$" to specify which version of the function you need.
--E 68
--       - 2967.0
--S 69 of 141
nagKelvinBer(60) :: Float
--R
--R   There are no library operations named nagKelvinBer
--R      Use HyperDoc Browse or issue
--R                            )what op nagKelvinBer
--R      to learn if there is any operation containing " nagKelvinBer " in
--R      its name.

--R
--RDaly Bug
--R   Cannot find a definition or applicable library operation named
--R      nagKelvinBer with argument type(s)
--R                               PositiveInteger
--R
--R      Perhaps you should use "@" to indicate the required return type,
--R      or "$" to specify which version of the function you need.
--E 69
--
-- ** ABNORMAL EXIT from NAG Library routine S19AAF: IFAIL =     1
-- ** NAG soft failure - control returned
--
--   Error signalled from user code:
--      An error was detected when calling the NAG Library routine
--      S19AAF. The error number (IFAIL value) is 1, please consult the
--      NAG manual via the Browser for diagnostic information.
--S 70 of 141
nagKelvinBer(-1) :: Float
--R
--R   There are no library operations named nagKelvinBer
--R      Use HyperDoc Browse or issue
--R                            )what op nagKelvinBer
--R      to learn if there is any operation containing " nagKelvinBer " in
--R      its name.
--R
--RDaly Bug
--R   Cannot find a definition or applicable library operation named
--R      nagKelvinBer with argument type(s)
--R                                   Integer
--R
--R      Perhaps you should use "@" to indicate the required return type,
--R      or "$" to specify which version of the function you need.
--E 70
--       0.9844

———-

nagKelvinBei : DF -¿ DF ; — * —

--S 71 of 141
nagKelvinBei(0.1) :: Float
--R
--R   There are no library operations named nagKelvinBei
--R      Use HyperDoc Browse or issue
--R                            )what op nagKelvinBei
--R      to learn if there is any operation containing " nagKelvinBei " in
--R      its name.

--R
--RDaly Bug
--R   Cannot find a definition or applicable library operation named
--R      nagKelvinBei with argument type(s)
--R                                    Float
--R
--R      Perhaps you should use "@" to indicate the required return type,
--R      or "$" to specify which version of the function you need.
--E 71
--       0.0025
--S 72 of 141
nagKelvinBei(1) :: Float
--R
--R   There are no library operations named nagKelvinBei
--R      Use HyperDoc Browse or issue
--R                            )what op nagKelvinBei
--R      to learn if there is any operation containing " nagKelvinBei " in
--R      its name.
--R
--RDaly Bug
--R   Cannot find a definition or applicable library operation named
--R      nagKelvinBei with argument type(s)
--R                               PositiveInteger
--R
--R      Perhaps you should use "@" to indicate the required return type,
--R      or "$" to specify which version of the function you need.
--E 72
--       0.2496
--S 73 of 141
nagKelvinBei(2.5) :: Float
--R
--R   There are no library operations named nagKelvinBei
--R      Use HyperDoc Browse or issue
--R                            )what op nagKelvinBei
--R      to learn if there is any operation containing " nagKelvinBei " in
--R      its name.
--R
--RDaly Bug
--R   Cannot find a definition or applicable library operation named
--R      nagKelvinBei with argument type(s)
--R                                    Float
--R
--R      Perhaps you should use "@" to indicate the required return type,
--R      or "$" to specify which version of the function you need.
--E 73
--       1.457
--S 74 of 141

nagKelvinBei(5) :: Float
--R
--R   There are no library operations named nagKelvinBei
--R      Use HyperDoc Browse or issue
--R                            )what op nagKelvinBei
--R      to learn if there is any operation containing " nagKelvinBei " in
--R      its name.
--R
--RDaly Bug
--R   Cannot find a definition or applicable library operation named
--R      nagKelvinBei with argument type(s)
--R                               PositiveInteger
--R
--R      Perhaps you should use "@" to indicate the required return type,
--R      or "$" to specify which version of the function you need.
--E 74
--       0.116
--S 75 of 141
nagKelvinBei(10) :: Float
--R
--R   There are no library operations named nagKelvinBei
--R      Use HyperDoc Browse or issue
--R                            )what op nagKelvinBei
--R      to learn if there is any operation containing " nagKelvinBei " in
--R      its name.
--R
--RDaly Bug
--R   Cannot find a definition or applicable library operation named
--R      nagKelvinBei with argument type(s)
--R                               PositiveInteger
--R
--R      Perhaps you should use "@" to indicate the required return type,
--R      or "$" to specify which version of the function you need.
--E 75
--       56.37
--S 76 of 141
nagKelvinBei(15) :: Float
--R
--R   There are no library operations named nagKelvinBei
--R      Use HyperDoc Browse or issue
--R                            )what op nagKelvinBei
--R      to learn if there is any operation containing " nagKelvinBei " in
--R      its name.
--R
--RDaly Bug
--R   Cannot find a definition or applicable library operation named
--R      nagKelvinBei with argument type(s)
--R                               PositiveInteger

--R
--R      Perhaps you should use "@" to indicate the required return type,
--R      or "$" to specify which version of the function you need.
--E 76
--       - 2953.0
--S 77 of 141
nagKelvinBei(60) :: Float
--R
--R   There are no library operations named nagKelvinBei
--R      Use HyperDoc Browse or issue
--R                            )what op nagKelvinBei
--R      to learn if there is any operation containing " nagKelvinBei " in
--R      its name.
--R
--RDaly Bug
--R   Cannot find a definition or applicable library operation named
--R      nagKelvinBei with argument type(s)
--R                               PositiveInteger
--R
--R      Perhaps you should use "@" to indicate the required return type,
--R      or "$" to specify which version of the function you need.
--E 77
--
-- ** ABNORMAL EXIT from NAG Library routine S19ABF: IFAIL =     1
-- ** NAG soft failure - control returned
--
--   Error signalled from user code:
--      An error was detected when calling the NAG Library routine
--      S19ABF. The error number (IFAIL value) is 1, please consult the
--      NAG manual via the Browser for diagnostic information.
--S 77a of 141
nagKelvinBei(-1) :: Float
--R
--R   There are no library operations named nagKelvinBei
--R      Use HyperDoc Browse or issue
--R                            )what op nagKelvinBei
--R      to learn if there is any operation containing " nagKelvinBei " in
--R      its name.
--R
--RDaly Bug
--R   Cannot find a definition or applicable library operation named
--R      nagKelvinBei with argument type(s)
--R                                   Integer
--R
--R      Perhaps you should use "@" to indicate the required return type,
--R      or "$" to specify which version of the function you need.
--E 77a
--       0.2496

———-

nagKelvinKer : DF -¿ DF ; — * —

--S 78 of 141
nagKelvinKer(0) :: Float
--R
--R   There are no library operations named nagKelvinKer
--R      Use HyperDoc Browse or issue
--R                            )what op nagKelvinKer
--R      to learn if there is any operation containing " nagKelvinKer " in
--R      its name.
--R
--RDaly Bug
--R   Cannot find a definition or applicable library operation named
--R      nagKelvinKer with argument type(s)
--R                             NonNegativeInteger
--R
--R      Perhaps you should use "@" to indicate the required return type,
--R      or "$" to specify which version of the function you need.
--E 78
--
-- ** ABNORMAL EXIT from NAG Library routine S19ACF: IFAIL =     2
-- ** NAG soft failure - control returned
--
--   Error signalled from user code:
--      An error was detected when calling the NAG Library routine
--      S19ACF. The error number (IFAIL value) is 2, please consult the
--      NAG manual via the Browser for diagnostic information.
--S 79 of 141
nagKelvinKer(0.1) :: Float
--R
--R   There are no library operations named nagKelvinKer
--R      Use HyperDoc Browse or issue
--R                            )what op nagKelvinKer
--R      to learn if there is any operation containing " nagKelvinKer " in
--R      its name.
--R
--RDaly Bug
--R   Cannot find a definition or applicable library operation named
--R      nagKelvinKer with argument type(s)
--R                                    Float
--R
--R      Perhaps you should use "@" to indicate the required return type,
--R      or "$" to specify which version of the function you need.
--E 79

--S 80 of 141
nagKelvinKer(1) :: Float
--R
--R   There are no library operations named nagKelvinKer
--R      Use HyperDoc Browse or issue
--R                            )what op nagKelvinKer
--R      to learn if there is any operation containing " nagKelvinKer " in
--R      its name.
--R
--RDaly Bug
--R   Cannot find a definition or applicable library operation named
--R      nagKelvinKer with argument type(s)
--R                               PositiveInteger
--R
--R      Perhaps you should use "@" to indicate the required return type,
--R      or "$" to specify which version of the function you need.
--E 80
--       0.2867
--S 81 of 141
nagKelvinKer(2.5) :: Float
--R
--R   There are no library operations named nagKelvinKer
--R      Use HyperDoc Browse or issue
--R                            )what op nagKelvinKer
--R      to learn if there is any operation containing " nagKelvinKer " in
--R      its name.
--R
--RDaly Bug
--R   Cannot find a definition or applicable library operation named
--R      nagKelvinKer with argument type(s)
--R                                    Float
--R
--R      Perhaps you should use "@" to indicate the required return type,
--R      or "$" to specify which version of the function you need.
--E 81
--       - 0.06969
--S 82 of 141
nagKelvinKer(5) :: Float
--R
--R   There are no library operations named nagKelvinKer
--R      Use HyperDoc Browse or issue
--R                            )what op nagKelvinKer
--R      to learn if there is any operation containing " nagKelvinKer " in
--R      its name.
--R
--RDaly Bug

--R   Cannot find a definition or applicable library operation named
--R      nagKelvinKer with argument type(s)
--R                               PositiveInteger
--R
--R      Perhaps you should use "@" to indicate the required return type,
--R      or "$" to specify which version of the function you need.
--E 82
--       - 0.01151
--S 83 of 141
nagKelvinKer(10) :: Float
--R
--R   There are no library operations named nagKelvinKer
--R      Use HyperDoc Browse or issue
--R                            )what op nagKelvinKer
--R      to learn if there is any operation containing " nagKelvinKer " in
--R      its name.
--R
--RDaly Bug
--R   Cannot find a definition or applicable library operation named
--R      nagKelvinKer with argument type(s)
--R                               PositiveInteger
--R
--R      Perhaps you should use "@" to indicate the required return type,
--R      or "$" to specify which version of the function you need.
--E 83
--       0.0001295
--S 84 of 141
nagKelvinKer(15) :: Float
--R
--R   There are no library operations named nagKelvinKer
--R      Use HyperDoc Browse or issue
--R                            )what op nagKelvinKer
--R      to learn if there is any operation containing " nagKelvinKer " in
--R      its name.
--R
--RDaly Bug
--R   Cannot find a definition or applicable library operation named
--R      nagKelvinKer with argument type(s)
--R                               PositiveInteger
--R
--R      Perhaps you should use "@" to indicate the required return type,
--R      or "$" to specify which version of the function you need.
--E 84
--       - 0.1514 E -7
--S 85 of 141
nagKelvinKer(1100) :: Float
--R

--R   There are no library operations named nagKelvinKer
--R      Use HyperDoc Browse or issue
--R                            )what op nagKelvinKer
--R      to learn if there is any operation containing " nagKelvinKer " in
--R      its name.
--R
--RDaly Bug
--R   Cannot find a definition or applicable library operation named
--R      nagKelvinKer with argument type(s)
--R                               PositiveInteger
--R
--R      Perhaps you should use "@" to indicate the required return type,
--R      or "$" to specify which version of the function you need.
--E 85
--
-- ** ABNORMAL EXIT from NAG Library routine S19ACF: IFAIL =     1
-- ** NAG soft failure - control returned
--
--   Error signalled from user code:
--      An error was detected when calling the NAG Library routine
--      S19ACF. The error number (IFAIL value) is 1, please consult the
--      NAG manual via the Browser for diagnostic information.
--S 86 of 141
nagKelvinKer(-1) :: Float
--R
--R   There are no library operations named nagKelvinKer
--R      Use HyperDoc Browse or issue
--R                            )what op nagKelvinKer
--R      to learn if there is any operation containing " nagKelvinKer " in
--R      its name.
--R
--RDaly Bug
--R   Cannot find a definition or applicable library operation named
--R      nagKelvinKer with argument type(s)
--R                                   Integer
--R
--R      Perhaps you should use "@" to indicate the required return type,
--R      or "$" to specify which version of the function you need.
--E 86
--
-- ** ABNORMAL EXIT from NAG Library routine S19ACF: IFAIL =     2
-- ** NAG soft failure - control returned
--
--   Error signalled from user code:
--      An error was detected when calling the NAG Library routine
--      S19ACF. The error number (IFAIL value) is 2, please consult the
--      NAG manual via the Browser for diagnostic information.

———-

nagKelvinKei : DF -¿ DF ; — * —

--S 87 of 141
nagKelvinKei(0) :: Float
--R
--R   There are no library operations named nagKelvinKei
--R      Use HyperDoc Browse or issue
--R                            )what op nagKelvinKei
--R      to learn if there is any operation containing " nagKelvinKei " in
--R      its name.
--R
--RDaly Bug
--R   Cannot find a definition or applicable library operation named
--R      nagKelvinKei with argument type(s)
--R                             NonNegativeInteger
--R
--R      Perhaps you should use "@" to indicate the required return type,
--R      or "$" to specify which version of the function you need.
--E 87
--       - 0.7854
--S 88 of 141
nagKelvinKei(0.1) :: Float
--R
--R   There are no library operations named nagKelvinKei
--R      Use HyperDoc Browse or issue
--R                            )what op nagKelvinKei
--R      to learn if there is any operation containing " nagKelvinKei " in
--R      its name.
--R
--RDaly Bug
--R   Cannot find a definition or applicable library operation named
--R      nagKelvinKei with argument type(s)
--R                                    Float
--R
--R      Perhaps you should use "@" to indicate the required return type,
--R      or "$" to specify which version of the function you need.
--E 88
--       - 0.7769
--S 89 of 141
nagKelvinKei(1) :: Float
--R
--R   There are no library operations named nagKelvinKei
--R      Use HyperDoc Browse or issue
--R                            )what op nagKelvinKei
--R      to learn if there is any operation containing " nagKelvinKei " in

--R      its name.
--R
--RDaly Bug
--R   Cannot find a definition or applicable library operation named
--R      nagKelvinKei with argument type(s)
--R                               PositiveInteger
--R
--R      Perhaps you should use "@" to indicate the required return type,
--R      or "$" to specify which version of the function you need.
--E 89
--       - 0.495
--S 90 of 141
nagKelvinKei(2.5) :: Float
--R
--R   There are no library operations named nagKelvinKei
--R      Use HyperDoc Browse or issue
--R                            )what op nagKelvinKei
--R      to learn if there is any operation containing " nagKelvinKei " in
--R      its name.
--R
--RDaly Bug
--R   Cannot find a definition or applicable library operation named
--R      nagKelvinKei with argument type(s)
--R                                    Float
--R
--R      Perhaps you should use "@" to indicate the required return type,
--R      or "$" to specify which version of the function you need.
--E 90
--       - 0.1107
--S 91 of 141
nagKelvinKei(5) :: Float
--R
--R   There are no library operations named nagKelvinKei
--R      Use HyperDoc Browse or issue
--R                            )what op nagKelvinKei
--R      to learn if there is any operation containing " nagKelvinKei " in
--R      its name.
--R
--RDaly Bug
--R   Cannot find a definition or applicable library operation named
--R      nagKelvinKei with argument type(s)
--R                               PositiveInteger
--R
--R      Perhaps you should use "@" to indicate the required return type,
--R      or "$" to specify which version of the function you need.
--E 91
--       0.01119

--S 92 of 141
nagKelvinKei(10) :: Float
--R
--R   There are no library operations named nagKelvinKei
--R      Use HyperDoc Browse or issue
--R                            )what op nagKelvinKei
--R      to learn if there is any operation containing " nagKelvinKei " in
--R      its name.
--R
--RDaly Bug
--R   Cannot find a definition or applicable library operation named
--R      nagKelvinKei with argument type(s)
--R                               PositiveInteger
--R
--R      Perhaps you should use "@" to indicate the required return type,
--R      or "$" to specify which version of the function you need.
--E 92
--       - 0.0003075
--S 93 of 141
nagKelvinKei(15) :: Float
--R
--R   There are no library operations named nagKelvinKei
--R      Use HyperDoc Browse or issue
--R                            )what op nagKelvinKei
--R      to learn if there is any operation containing " nagKelvinKei " in
--R      its name.
--R
--RDaly Bug
--R   Cannot find a definition or applicable library operation named
--R      nagKelvinKei with argument type(s)
--R                               PositiveInteger
--R
--R      Perhaps you should use "@" to indicate the required return type,
--R      or "$" to specify which version of the function you need.
--E 93
--       0.000007963
--S 94 of 141
nagKelvinKei(1100) :: Float
--R
--R   There are no library operations named nagKelvinKei
--R      Use HyperDoc Browse or issue
--R                            )what op nagKelvinKei
--R      to learn if there is any operation containing " nagKelvinKei " in
--R      its name.
--R
--RDaly Bug
--R   Cannot find a definition or applicable library operation named
--R      nagKelvinKei with argument type(s)

--R                               PositiveInteger
--R
--R      Perhaps you should use "@" to indicate the required return type,
--R      or "$" to specify which version of the function you need.
--E 94
--
-- ** ABNORMAL EXIT from NAG Library routine S19ADF: IFAIL =     1
-- ** NAG soft failure - control returned
--
--   Error signalled from user code:
--      An error was detected when calling the NAG Library routine
--      S19ADF. The error number (IFAIL value) is 1, please consult the
--      NAG manual via the Browser for diagnostic information.
--S 95 of 141
nagKelvinKei(-1) :: Float
--R
--R   There are no library operations named nagKelvinKei
--R      Use HyperDoc Browse or issue
--R                            )what op nagKelvinKei
--R      to learn if there is any operation containing " nagKelvinKei " in
--R      its name.
--R
--RDaly Bug
--R   Cannot find a definition or applicable library operation named
--R      nagKelvinKei with argument type(s)
--R                                   Integer
--R
--R      Perhaps you should use "@" to indicate the required return type,
--R      or "$" to specify which version of the function you need.
--E 95
--
-- ** ABNORMAL EXIT from NAG Library routine S19ADF: IFAIL =     2
-- ** NAG soft failure - control returned
--
--   Error signalled from user code:
--      An error was detected when calling the NAG Library routine
--      S19ADF. The error number (IFAIL value) is 2, please consult the
--      NAG manual via the Browser for diagnostic information.

———-

nagFresnelS : DF -¿ DF ; — * —

--S 96 of 141
nagFresnelS(0) :: Float
--R
--R   There are no library operations named nagFresnelS

--R      Use HyperDoc Browse or issue
--R                            )what op nagFresnelS
--R      to learn if there is any operation containing " nagFresnelS " in
--R      its name.
--R
--RDaly Bug
--R   Cannot find a definition or applicable library operation named
--R      nagFresnelS with argument type(s)
--R                             NonNegativeInteger
--R
--R      Perhaps you should use "@" to indicate the required return type,
--R      or "$" to specify which version of the function you need.
--E 96
--       0.0
--S 97 of 141
nagFresnelS(0.5) :: Float
--R
--R   There are no library operations named nagFresnelS
--R      Use HyperDoc Browse or issue
--R                            )what op nagFresnelS
--R      to learn if there is any operation containing " nagFresnelS " in
--R      its name.
--R
--RDaly Bug
--R   Cannot find a definition or applicable library operation named
--R      nagFresnelS with argument type(s)
--R                                    Float
--R
--R      Perhaps you should use "@" to indicate the required return type,
--R      or "$" to specify which version of the function you need.
--E 97
--       0.06473
--S 98 of 141
nagFresnelS(1) :: Float
--R
--R   There are no library operations named nagFresnelS
--R      Use HyperDoc Browse or issue
--R                            )what op nagFresnelS
--R      to learn if there is any operation containing " nagFresnelS " in
--R      its name.
--R
--RDaly Bug
--R   Cannot find a definition or applicable library operation named
--R      nagFresnelS with argument type(s)
--R                               PositiveInteger
--R
--R      Perhaps you should use "@" to indicate the required return type,
--R      or "$" to specify which version of the function you need.

--E 98
--       0.4383
--S 99 of 141
nagFresnelS(2) :: Float
--R
--R   There are no library operations named nagFresnelS
--R      Use HyperDoc Browse or issue
--R                            )what op nagFresnelS
--R      to learn if there is any operation containing " nagFresnelS " in
--R      its name.
--R
--RDaly Bug
--R   Cannot find a definition or applicable library operation named
--R      nagFresnelS with argument type(s)
--R                               PositiveInteger
--R
--R      Perhaps you should use "@" to indicate the required return type,
--R      or "$" to specify which version of the function you need.
--E 99
--       0.3434
--S 100 of 141
nagFresnelS(4) :: Float
--R
--R   There are no library operations named nagFresnelS
--R      Use HyperDoc Browse or issue
--R                            )what op nagFresnelS
--R      to learn if there is any operation containing " nagFresnelS " in
--R      its name.
--R
--RDaly Bug
--R   Cannot find a definition or applicable library operation named
--R      nagFresnelS with argument type(s)
--R                               PositiveInteger
--R
--R      Perhaps you should use "@" to indicate the required return type,
--R      or "$" to specify which version of the function you need.
--E 100
--       0.4205
--S 101 of 141
nagFresnelS(5) :: Float
--R
--R   There are no library operations named nagFresnelS
--R      Use HyperDoc Browse or issue
--R                            )what op nagFresnelS
--R      to learn if there is any operation containing " nagFresnelS " in
--R      its name.
--R

--RDaly Bug
--R   Cannot find a definition or applicable library operation named
--R      nagFresnelS with argument type(s)
--R                               PositiveInteger
--R
--R      Perhaps you should use "@" to indicate the required return type,
--R      or "$" to specify which version of the function you need.
--E 101
--       0.4992
--S 102 of 141
nagFresnelS(6) :: Float
--R
--R   There are no library operations named nagFresnelS
--R      Use HyperDoc Browse or issue
--R                            )what op nagFresnelS
--R      to learn if there is any operation containing " nagFresnelS " in
--R      its name.
--R
--RDaly Bug
--R   Cannot find a definition or applicable library operation named
--R      nagFresnelS with argument type(s)
--R                               PositiveInteger
--R
--R      Perhaps you should use "@" to indicate the required return type,
--R      or "$" to specify which version of the function you need.
--E 102
--       0.447
--S 103 of 141
nagFresnelS(8) :: Float
--R
--R   There are no library operations named nagFresnelS
--R      Use HyperDoc Browse or issue
--R                            )what op nagFresnelS
--R      to learn if there is any operation containing " nagFresnelS " in
--R      its name.
--R
--RDaly Bug
--R   Cannot find a definition or applicable library operation named
--R      nagFresnelS with argument type(s)
--R                               PositiveInteger
--R
--R      Perhaps you should use "@" to indicate the required return type,
--R      or "$" to specify which version of the function you need.
--E 103
--       0.4602
--S 104 of 141
nagFresnelS(10) :: Float

--R
--R   There are no library operations named nagFresnelS
--R      Use HyperDoc Browse or issue
--R                            )what op nagFresnelS
--R      to learn if there is any operation containing " nagFresnelS " in
--R      its name.
--R
--RDaly Bug
--R   Cannot find a definition or applicable library operation named
--R      nagFresnelS with argument type(s)
--R                               PositiveInteger
--R
--R      Perhaps you should use "@" to indicate the required return type,
--R      or "$" to specify which version of the function you need.
--E 104
--       0.4682
--S 105 of 141
nagFresnelS(-1) :: Float
--R
--R   There are no library operations named nagFresnelS
--R      Use HyperDoc Browse or issue
--R                            )what op nagFresnelS
--R      to learn if there is any operation containing " nagFresnelS " in
--R      its name.
--R
--RDaly Bug
--R   Cannot find a definition or applicable library operation named
--R      nagFresnelS with argument type(s)
--R                                   Integer
--R
--R      Perhaps you should use "@" to indicate the required return type,
--R      or "$" to specify which version of the function you need.
--E 105
--       - 0.4383
--S 106 of 141
nagFresnelS(1000) :: Float
--R
--R   There are no library operations named nagFresnelS
--R      Use HyperDoc Browse or issue
--R                            )what op nagFresnelS
--R      to learn if there is any operation containing " nagFresnelS " in
--R      its name.
--R
--RDaly Bug
--R   Cannot find a definition or applicable library operation named
--R      nagFresnelS with argument type(s)
--R                               PositiveInteger
--R

--R      Perhaps you should use "@" to indicate the required return type,
--R      or "$" to specify which version of the function you need.
--E 106
--       0.4997

———-

nagFresnelC : DF -¿ DF ; — * —

--S 107 of 141
nagFresnelC(0) :: Float
--R
--R   There are no library operations named nagFresnelC
--R      Use HyperDoc Browse or issue
--R                            )what op nagFresnelC
--R      to learn if there is any operation containing " nagFresnelC " in
--R      its name.
--R
--RDaly Bug
--R   Cannot find a definition or applicable library operation named
--R      nagFresnelC with argument type(s)
--R                             NonNegativeInteger
--R
--R      Perhaps you should use "@" to indicate the required return type,
--R      or "$" to specify which version of the function you need.
--E 107
--       0.0
--S 108 of 141
nagFresnelC(0.5) :: Float
--R
--R   There are no library operations named nagFresnelC
--R      Use HyperDoc Browse or issue
--R                            )what op nagFresnelC
--R      to learn if there is any operation containing " nagFresnelC " in
--R      its name.
--R
--RDaly Bug
--R   Cannot find a definition or applicable library operation named
--R      nagFresnelC with argument type(s)
--R                                    Float
--R
--R      Perhaps you should use "@" to indicate the required return type,
--R      or "$" to specify which version of the function you need.
--E 108
--       0.4923
--S 109 of 141

nagFresnelC(1) :: Float
--R
--R   There are no library operations named nagFresnelC
--R      Use HyperDoc Browse or issue
--R                            )what op nagFresnelC
--R      to learn if there is any operation containing " nagFresnelC " in
--R      its name.
--R
--RDaly Bug
--R   Cannot find a definition or applicable library operation named
--R      nagFresnelC with argument type(s)
--R                               PositiveInteger
--R
--R      Perhaps you should use "@" to indicate the required return type,
--R      or "$" to specify which version of the function you need.
--E 109
--       0.7799
--S 110 of 141
nagFresnelC(2) :: Float
--R
--R   There are no library operations named nagFresnelC
--R      Use HyperDoc Browse or issue
--R                            )what op nagFresnelC
--R      to learn if there is any operation containing " nagFresnelC " in
--R      its name.
--R
--RDaly Bug
--R   Cannot find a definition or applicable library operation named
--R      nagFresnelC with argument type(s)
--R                               PositiveInteger
--R
--R      Perhaps you should use "@" to indicate the required return type,
--R      or "$" to specify which version of the function you need.
--E 110
--       0.4883
--S 111 of 141
nagFresnelC(4) :: Float
--R
--R   There are no library operations named nagFresnelC
--R      Use HyperDoc Browse or issue
--R                            )what op nagFresnelC
--R      to learn if there is any operation containing " nagFresnelC " in
--R      its name.
--R
--RDaly Bug
--R   Cannot find a definition or applicable library operation named
--R      nagFresnelC with argument type(s)
--R                               PositiveInteger

--R
--R      Perhaps you should use "@" to indicate the required return type,
--R      or "$" to specify which version of the function you need.
--E 111
--       0.4984
--S 112 of 141
nagFresnelC(5) :: Float
--R
--R   There are no library operations named nagFresnelC
--R      Use HyperDoc Browse or issue
--R                            )what op nagFresnelC
--R      to learn if there is any operation containing " nagFresnelC " in
--R      its name.
--R
--RDaly Bug
--R   Cannot find a definition or applicable library operation named
--R      nagFresnelC with argument type(s)
--R                               PositiveInteger
--R
--R      Perhaps you should use "@" to indicate the required return type,
--R      or "$" to specify which version of the function you need.
--E 112
--       0.5636
--S 113 of 141
nagFresnelC(6) :: Float
--R
--R   There are no library operations named nagFresnelC
--R      Use HyperDoc Browse or issue
--R                            )what op nagFresnelC
--R      to learn if there is any operation containing " nagFresnelC " in
--R      its name.
--R
--RDaly Bug
--R   Cannot find a definition or applicable library operation named
--R      nagFresnelC with argument type(s)
--R                               PositiveInteger
--R
--R      Perhaps you should use "@" to indicate the required return type,
--R      or "$" to specify which version of the function you need.
--E 113
--       0.4995
--S 114 of 141
nagFresnelC(8) :: Float
--R
--R   There are no library operations named nagFresnelC
--R      Use HyperDoc Browse or issue
--R                            )what op nagFresnelC

--R      to learn if there is any operation containing " nagFresnelC " in
--R      its name.
--R
--RDaly Bug
--R   Cannot find a definition or applicable library operation named
--R      nagFresnelC with argument type(s)
--R                               PositiveInteger
--R
--R      Perhaps you should use "@" to indicate the required return type,
--R      or "$" to specify which version of the function you need.
--E 114
--       0.4998
--S 115 of 141
nagFresnelC(10) :: Float
--R
--R   There are no library operations named nagFresnelC
--R      Use HyperDoc Browse or issue
--R                            )what op nagFresnelC
--R      to learn if there is any operation containing " nagFresnelC " in
--R      its name.
--R
--RDaly Bug
--R   Cannot find a definition or applicable library operation named
--R      nagFresnelC with argument type(s)
--R                               PositiveInteger
--R
--R      Perhaps you should use "@" to indicate the required return type,
--R      or "$" to specify which version of the function you need.
--E 115
--       0.4999
--S 116 of 141
nagFresnelC(-1) :: Float
--R
--R   There are no library operations named nagFresnelC
--R      Use HyperDoc Browse or issue
--R                            )what op nagFresnelC
--R      to learn if there is any operation containing " nagFresnelC " in
--R      its name.
--R
--RDaly Bug
--R   Cannot find a definition or applicable library operation named
--R      nagFresnelC with argument type(s)
--R                                   Integer
--R
--R      Perhaps you should use "@" to indicate the required return type,
--R      or "$" to specify which version of the function you need.
--E 116
--       - 0.7799

--S 117 of 141
nagFresnelC(1000) :: Float
--R
--R   There are no library operations named nagFresnelC
--R      Use HyperDoc Browse or issue
--R                            )what op nagFresnelC
--R      to learn if there is any operation containing " nagFresnelC " in
--R      its name.
--R
--RDaly Bug
--R   Cannot find a definition or applicable library operation named
--R      nagFresnelC with argument type(s)
--R                               PositiveInteger
--R
--R      Perhaps you should use "@" to indicate the required return type,
--R      or "$" to specify which version of the function you need.
--E 117
--       0.5

———-

– nagEllipticIntegralRC : (DF, DF) -¿ DF ; — * —

--S 118 of 141
nagEllipticIntegralRC(0.5,1) :: Float
--R
--R   There are no library operations named nagEllipticIntegralRC
--R      Use HyperDoc Browse or issue
--R                       )what op nagEllipticIntegralRC
--R      to learn if there is any operation containing "
--R      nagEllipticIntegralRC " in its name.
--R
--RDaly Bug
--R   Cannot find a definition or applicable library operation named
--R      nagEllipticIntegralRC with argument type(s)
--R                                    Float
--R                               PositiveInteger
--R
--R      Perhaps you should use "@" to indicate the required return type,
--R      or "$" to specify which version of the function you need.
--E 118
--       1.111
--S 119 of 141
nagEllipticIntegralRC(1,1) :: Float
--R
--R   There are no library operations named nagEllipticIntegralRC

--R      Use HyperDoc Browse or issue
--R                       )what op nagEllipticIntegralRC
--R      to learn if there is any operation containing "
--R      nagEllipticIntegralRC " in its name.
--R
--RDaly Bug
--R   Cannot find a definition or applicable library operation named
--R      nagEllipticIntegralRC with argument type(s)
--R                               PositiveInteger
--R                               PositiveInteger
--R
--R      Perhaps you should use "@" to indicate the required return type,
--R      or "$" to specify which version of the function you need.
--E 119
--       1.0
--S 120 of 141
nagEllipticIntegralRC(1.5,1) :: Float
--R
--R   There are no library operations named nagEllipticIntegralRC
--R      Use HyperDoc Browse or issue
--R                       )what op nagEllipticIntegralRC
--R      to learn if there is any operation containing "
--R      nagEllipticIntegralRC " in its name.
--R
--RDaly Bug
--R   Cannot find a definition or applicable library operation named
--R      nagEllipticIntegralRC with argument type(s)
--R                                    Float
--R                               PositiveInteger
--R
--R      Perhaps you should use "@" to indicate the required return type,
--R      or "$" to specify which version of the function you need.
--E 120
--       0.9312

———-

nagEllipticIntegralRD : (DF, DF, DF) -¿ DF ; — * —

--S 121 of 141
nagEllipticIntegralRD(0.5,0.5,1) :: Float
--R
--R   There are no library operations named nagEllipticIntegralRD
--R      Use HyperDoc Browse or issue
--R                       )what op nagEllipticIntegralRD
--R      to learn if there is any operation containing "
--R      nagEllipticIntegralRD " in its name.
--R

--RDaly Bug
--R   Cannot find a definition or applicable library operation named
--R      nagEllipticIntegralRD with argument type(s)
--R                                    Float
--R                                    Float
--R                               PositiveInteger
--R
--R      Perhaps you should use "@" to indicate the required return type,
--R      or "$" to specify which version of the function you need.
--E 121
--       1.479
--S 122 of 141
nagEllipticIntegralRD(0.5,1,1) :: Float
--R
--R   There are no library operations named nagEllipticIntegralRD
--R      Use HyperDoc Browse or issue
--R                       )what op nagEllipticIntegralRD
--R      to learn if there is any operation containing "
--R      nagEllipticIntegralRD " in its name.
--R
--RDaly Bug
--R   Cannot find a definition or applicable library operation named
--R      nagEllipticIntegralRD with argument type(s)
--R                                    Float
--R                               PositiveInteger
--R                               PositiveInteger
--R
--R      Perhaps you should use "@" to indicate the required return type,
--R      or "$" to specify which version of the function you need.
--E 122
--       1.211
--S 123 of 141
nagEllipticIntegralRD(0.5,1.5,1) :: Float
--R
--R   There are no library operations named nagEllipticIntegralRD
--R      Use HyperDoc Browse or issue
--R                       )what op nagEllipticIntegralRD
--R      to learn if there is any operation containing "
--R      nagEllipticIntegralRD " in its name.
--R
--RDaly Bug
--R   Cannot find a definition or applicable library operation named
--R      nagEllipticIntegralRD with argument type(s)
--R                                    Float
--R                                    Float
--R                               PositiveInteger
--R
--R      Perhaps you should use "@" to indicate the required return type,

--R      or "$" to specify which version of the function you need.
--E 123
--       1.061
--S 124 of 141
nagEllipticIntegralRD(1,1,1) :: Float
--R
--R   There are no library operations named nagEllipticIntegralRD
--R      Use HyperDoc Browse or issue
--R                       )what op nagEllipticIntegralRD
--R      to learn if there is any operation containing "
--R      nagEllipticIntegralRD " in its name.
--R
--RDaly Bug
--R   Cannot find a definition or applicable library operation named
--R      nagEllipticIntegralRD with argument type(s)
--R                               PositiveInteger
--R                               PositiveInteger
--R                               PositiveInteger
--R
--R      Perhaps you should use "@" to indicate the required return type,
--R      or "$" to specify which version of the function you need.
--E 124
--       1.0
--S 125 of 141
nagEllipticIntegralRD(1,1.5,1) :: Float
--R
--R   There are no library operations named nagEllipticIntegralRD
--R      Use HyperDoc Browse or issue
--R                       )what op nagEllipticIntegralRD
--R      to learn if there is any operation containing "
--R      nagEllipticIntegralRD " in its name.
--R
--RDaly Bug
--R   Cannot find a definition or applicable library operation named
--R      nagEllipticIntegralRD with argument type(s)
--R                               PositiveInteger
--R                                    Float
--R                               PositiveInteger
--R
--R      Perhaps you should use "@" to indicate the required return type,
--R      or "$" to specify which version of the function you need.
--E 125
--       0.8805
--S 126 of 141
nagEllipticIntegralRD(1.5,1.5,1) :: Float
--R
--R   There are no library operations named nagEllipticIntegralRD

--R      Use HyperDoc Browse or issue
--R                       )what op nagEllipticIntegralRD
--R      to learn if there is any operation containing "
--R      nagEllipticIntegralRD " in its name.
--R
--RDaly Bug
--R   Cannot find a definition or applicable library operation named
--R      nagEllipticIntegralRD with argument type(s)
--R                                    Float
--R                                    Float
--R                               PositiveInteger
--R
--R      Perhaps you should use "@" to indicate the required return type,
--R      or "$" to specify which version of the function you need.
--E 126
--       0.7775

———-

nagEllipticIntegralRF : (DF, DF, DF) -¿ DF ; — * —

--S 127 of 141
nagEllipticIntegralRF(0.5,1,1.5) :: Float
--R
--R   There are no library operations named nagEllipticIntegralRF
--R      Use HyperDoc Browse or issue
--R                       )what op nagEllipticIntegralRF
--R      to learn if there is any operation containing "
--R      nagEllipticIntegralRF " in its name.
--R
--RDaly Bug
--R   Cannot find a definition or applicable library operation named
--R      nagEllipticIntegralRF with argument type(s)
--R                                    Float
--R                               PositiveInteger
--R                                    Float
--R
--R      Perhaps you should use "@" to indicate the required return type,
--R      or "$" to specify which version of the function you need.
--E 127
--       1.028
--S 128 of 141
nagEllipticIntegralRF(1,1.5,2) :: Float
--R
--R   There are no library operations named nagEllipticIntegralRF
--R      Use HyperDoc Browse or issue
--R                       )what op nagEllipticIntegralRF
--R      to learn if there is any operation containing "

--R      nagEllipticIntegralRF " in its name.
--R
--RDaly Bug
--R   Cannot find a definition or applicable library operation named
--R      nagEllipticIntegralRF with argument type(s)
--R                               PositiveInteger
--R                                    Float
--R                               PositiveInteger
--R
--R      Perhaps you should use "@" to indicate the required return type,
--R      or "$" to specify which version of the function you need.
--E 128
--       0.826
--S 129 of 141
nagEllipticIntegralRF(1.5,2,2.5) :: Float
--R
--R   There are no library operations named nagEllipticIntegralRF
--R      Use HyperDoc Browse or issue
--R                       )what op nagEllipticIntegralRF
--R      to learn if there is any operation containing "
--R      nagEllipticIntegralRF " in its name.
--R
--RDaly Bug
--R   Cannot find a definition or applicable library operation named
--R      nagEllipticIntegralRF with argument type(s)
--R                                    Float
--R                               PositiveInteger
--R                                    Float
--R
--R      Perhaps you should use "@" to indicate the required return type,
--R      or "$" to specify which version of the function you need.
--E 129
--       0.7116

———-

nagEllipticIntegralRJ : (DF, DF, DF, DF) -¿ DF ; — * —

--S 130 of 141
nagEllipticIntegralRJ(0.5,0.5,0.5,2) :: Float
--R
--R   There are no library operations named nagEllipticIntegralRJ
--R      Use HyperDoc Browse or issue
--R                       )what op nagEllipticIntegralRJ
--R      to learn if there is any operation containing "
--R      nagEllipticIntegralRJ " in its name.
--R
--RDaly Bug

--R   Cannot find a definition or applicable library operation named
--R      nagEllipticIntegralRJ with argument type(s)
--R                                    Float
--R                                    Float
--R                                    Float
--R                               PositiveInteger
--R
--R      Perhaps you should use "@" to indicate the required return type,
--R      or "$" to specify which version of the function you need.
--E 130
--       1.118
--S 131 of 141
nagEllipticIntegralRJ(0.5,0.5,1,2) :: Float
--R
--R   There are no library operations named nagEllipticIntegralRJ
--R      Use HyperDoc Browse or issue
--R                       )what op nagEllipticIntegralRJ
--R      to learn if there is any operation containing "
--R      nagEllipticIntegralRJ " in its name.
--R
--RDaly Bug
--R   Cannot find a definition or applicable library operation named
--R      nagEllipticIntegralRJ with argument type(s)
--R                                    Float
--R                                    Float
--R                               PositiveInteger
--R                               PositiveInteger
--R
--R      Perhaps you should use "@" to indicate the required return type,
--R      or "$" to specify which version of the function you need.
--E 131
--       0.9221
--S 132 of 141
nagEllipticIntegralRJ(0.5,0.5,1.5,2) :: Float
--R
--R   There are no library operations named nagEllipticIntegralRJ
--R      Use HyperDoc Browse or issue
--R                       )what op nagEllipticIntegralRJ
--R      to learn if there is any operation containing "
--R      nagEllipticIntegralRJ " in its name.
--R
--RDaly Bug
--R   Cannot find a definition or applicable library operation named
--R      nagEllipticIntegralRJ with argument type(s)
--R                                    Float
--R                                    Float
--R                                    Float
--R                               PositiveInteger

--R
--R      Perhaps you should use "@" to indicate the required return type,
--R      or "$" to specify which version of the function you need.
--E 132
--       0.8115
--S 133 of 141
nagEllipticIntegralRJ(0.5,1,1,2) :: Float
--R
--R   There are no library operations named nagEllipticIntegralRJ
--R      Use HyperDoc Browse or issue
--R                       )what op nagEllipticIntegralRJ
--R      to learn if there is any operation containing "
--R      nagEllipticIntegralRJ " in its name.
--R
--RDaly Bug
--R   Cannot find a definition or applicable library operation named
--R      nagEllipticIntegralRJ with argument type(s)
--R                                    Float
--R                               PositiveInteger
--R                               PositiveInteger
--R                               PositiveInteger
--R
--R      Perhaps you should use "@" to indicate the required return type,
--R      or "$" to specify which version of the function you need.
--E 133
--       0.7671
--S 134 of 141
nagEllipticIntegralRJ(0.5,1,1.5,2) :: Float
--R
--R   There are no library operations named nagEllipticIntegralRJ
--R      Use HyperDoc Browse or issue
--R                       )what op nagEllipticIntegralRJ
--R      to learn if there is any operation containing "
--R      nagEllipticIntegralRJ " in its name.
--R
--RDaly Bug
--R   Cannot find a definition or applicable library operation named
--R      nagEllipticIntegralRJ with argument type(s)
--R                                    Float
--R                               PositiveInteger
--R                                    Float
--R                               PositiveInteger
--R
--R      Perhaps you should use "@" to indicate the required return type,
--R      or "$" to specify which version of the function you need.
--E 134
--       0.6784

--S 135 of 141
nagEllipticIntegralRJ(0.5,1.5,1.5,2) :: Float
--R
--R   There are no library operations named nagEllipticIntegralRJ
--R      Use HyperDoc Browse or issue
--R                       )what op nagEllipticIntegralRJ
--R      to learn if there is any operation containing "
--R      nagEllipticIntegralRJ " in its name.
--R
--RDaly Bug
--R   Cannot find a definition or applicable library operation named
--R      nagEllipticIntegralRJ with argument type(s)
--R                                    Float
--R                                    Float
--R                                    Float
--R                               PositiveInteger
--R
--R      Perhaps you should use "@" to indicate the required return type,
--R      or "$" to specify which version of the function you need.
--E 135
--       0.6017
--S 136 of 141
nagEllipticIntegralRJ(1,1,1,2) :: Float
--R
--R   There are no library operations named nagEllipticIntegralRJ
--R      Use HyperDoc Browse or issue
--R                       )what op nagEllipticIntegralRJ
--R      to learn if there is any operation containing "
--R      nagEllipticIntegralRJ " in its name.
--R
--RDaly Bug
--R   Cannot find a definition or applicable library operation named
--R      nagEllipticIntegralRJ with argument type(s)
--R                               PositiveInteger
--R                               PositiveInteger
--R                               PositiveInteger
--R                               PositiveInteger
--R
--R      Perhaps you should use "@" to indicate the required return type,
--R      or "$" to specify which version of the function you need.
--E 136
--       0.6438
--S 137 of 141
nagEllipticIntegralRJ(1,1,1.5,2) :: Float
--R
--R   There are no library operations named nagEllipticIntegralRJ
--R      Use HyperDoc Browse or issue
--R                       )what op nagEllipticIntegralRJ

--R      to learn if there is any operation containing "
--R      nagEllipticIntegralRJ " in its name.
--R
--RDaly Bug
--R   Cannot find a definition or applicable library operation named
--R      nagEllipticIntegralRJ with argument type(s)
--R                               PositiveInteger
--R                               PositiveInteger
--R                                    Float
--R                               PositiveInteger
--R
--R      Perhaps you should use "@" to indicate the required return type,
--R      or "$" to specify which version of the function you need.
--E 137
--       0.5722
--S 138 of 141
nagEllipticIntegralRJ(1,1.5,1.5,2) :: Float
--R
--R   There are no library operations named nagEllipticIntegralRJ
--R      Use HyperDoc Browse or issue
--R                       )what op nagEllipticIntegralRJ
--R      to learn if there is any operation containing "
--R      nagEllipticIntegralRJ " in its name.
--R
--RDaly Bug
--R   Cannot find a definition or applicable library operation named
--R      nagEllipticIntegralRJ with argument type(s)
--R                               PositiveInteger
--R                                    Float
--R                                    Float
--R                               PositiveInteger
--R
--R      Perhaps you should use "@" to indicate the required return type,
--R      or "$" to specify which version of the function you need.
--E 138
--       0.5101
--S 139 of 141
nagEllipticIntegralRJ(1.5,1.5,1.5,2) :: Float
--R
--R   There are no library operations named nagEllipticIntegralRJ
--R      Use HyperDoc Browse or issue
--R                       )what op nagEllipticIntegralRJ
--R      to learn if there is any operation containing "
--R      nagEllipticIntegralRJ " in its name.
--R
--RDaly Bug
--R   Cannot find a definition or applicable library operation named
--R      nagEllipticIntegralRJ with argument type(s)

--R                                    Float
--R                                    Float
--R                                    Float
--R                               PositiveInteger
--R
--R      Perhaps you should use "@" to indicate the required return type,
--R      or "$" to specify which version of the function you need.
--E 139
--       0.4561
--S 140 of 141
outputGeneral()
--R
--R                                                                   Type: Void
--E 140
--S 141 of 141
output "End of tests"
--R
--R   End of tests
--R                                                                   Type: Void
--E 141
)spool
)lisp (bye)

———-

References^{[1] nothing}

63